Sales ​$50,063,085    Variable costs (28,483,000) Revenue before fixed costs ​$21,580,085    Fixed costs (15,457,000) EBIT ​$6,123,085    Interest...

Sales

​$50,063,085

  

Variable costs

(28,483,000)

Revenue before fixed costs

​$21,580,085

  

Fixed costs

(15,457,000)

EBIT

​$6,123,085

  

Interest expense

(1,337,331)

Earnings before taxes

​$4,785,754

  

Taxes at %50%

(2,392,877)

Net income

   $2,392,877  

(

Break-even

analysis​)

You have developed the income statement in the popup​ window,

​, for the Hugo Boss Corporation. It represents the most recent​ year's operations, which ended yesterday. Your supervisor in the​ controller's office has just handed you a memorandum asking for written responses to the following​ questions:

a. What is the​ firm's break-even point in sales​ dollars?

b. If sales should increase by

20

​percent, by what percent would earnings before taxes​ (and net​ income) increase?

a. What is the​ firm's break-even point in sales​ dollars?

​$   

​(Round to the nearest​ dollar.)
Answer #1

1.
Firm's break-even point in sales dollars=(15457000+1337331)/(1-28483000/50063085)=38960737.1969

2.
% increase in EBIT=(50063085*1.1-28483000*1.1-15457000)/6123085-1=35.244%

% increase in Net Income=(50063085*1.1-28483000*1.1-15457000-1337331)*(1-50%)/2392877-1=45.092%